# Algorithmic Misjudgement in Google Search Results: Evidence from Auditing the US Online Electoral Information Environment

### v1

@@ Abstract.
Google Search is an important way that people seek information about politics (Dutton et al., 2017), and Google states that it is “committed to providing timely and authoritative information on Google Search to help voters understand, navigate, and participate in democratic processes”. [^1] https://elections.google/civics-in-search/ In this paper, we interrogate the extent to which government-maintained web domains are represented in the online environment of electoral information of the 2022 US midterm elections, as captured through Google Search results in 3.45 million SERPs for 786 locations across the United States between October and November 2022. Although we find that almost 40% of organic results are contributed by the 40% of government domains, this proportional equilibrium hides the fact that most results either belong to a small number of popular domains or are mistargeted (at a rate of 71.18%) with respect to the location of the search. We consider the frequent omission and mistargeting of non-federal websites a form of algorithmic misjudgement that contributes to civic harm, by obscuring the important role that these institutions play in the election information environment.
@@
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: ACM Conference on Fairness, Accountability, and Transparency; June 03–06, 2024; Rio de Janeiro, Brazilisbn: 978-1-4503-XXXX-X/18/06
@@ 1. Introduction
Electoral denialism (the belief that elections are rigged or that electoral fraud is rampant) in the United States (but also other countries, such as Brasil and Peru) is emerging as a significant threat to democracy (Cameron, 2023), and it has roots, among others in racial resentment and rising conspiracism (Stewart III, 2023). Since the early 2000s, internet was regarded as a medium “uniquely receptive to conspiracism” (James, 2000), but in the past two decades, there is little doubt that it has contributed significantly to conspiracism going mainstream (Rosenblum and Muirhead, 2019). Following the 2020 US Presidential Election, “election rigging” conspiracies that came to be labeled as “The Big Lie” led among others to the violent attack to the US Capitol on January 6 2021 (Fahey, 2023). While beliefs about electoral fraud are somewhat persistent (Enders et al., 2021), the 2020 electoral denialism movement was different because it was embraced by partisan state legislatures, as reported by the Brennan Center for Justice. [^2] https://www.brennancenter.org/our-work/research-reports/voting-laws-roundup-june-2023 As a result, a large number of changes to electoral laws (some of them restricting and some of them expanding voting rights), became operational ahead of the 2022 US midterm election. [^3] https://www.reuters.com/legal/how-new-us-laws-could-trip-up-voters-midterm-elections-2022-11-01/
@@
Likewise, search engines also position themselves as trustworthy intermediaries for seeking election information. Google states that it is “committed to providing timely and authoritative information on Google Search to help voters understand, navigate, and participate in democratic processes.” [^6] https://elections.google/civics-in-search/ In an article about their work to “support” the 2022 US midterm elections, Google’s Vice President for Trust and Safety writes, “Our work is centered around connecting voters to the latest election information …” [^7] https://blog.google/outreach-initiatives/civics/our-ongoing-work-to-support-the-2022-us-midterm-elections/ and highlights “features that show data from nonpartisan organizations” that work to “connect voters with accurate information about voter registration and how to vote,” referring to Google Search election information panels that contain aggregated information from Democracy Works [^8] Democracy Works describes itself as “a nonpartisan, nonprofit organization that collaborates with election officials, leading tech platforms, and world-class partners to drive voter access and participation”. Source: https://www.democracy.works/about. about where and how to vote.
@@
Study Motivation: Given this background, three assumptions motivate the current study: (1) voting-eligible US citizens who search about electoral information using Google’s search engine have the right to receive accurate and relevant search results, regardless of where they live, 2) similar to Ballatore et al. (Andrea Ballatore and Sen, 2017), we acknowledge a spatial dimension to searching and search results, in which users searching for information related to elections and voting deserve results relevant to their geographic area, and 3) given the size of the United States and the varying voting laws per state (most of them updated ahead of the 2022 elections), receiving information from authoritative, locality-specific sources is a reasonable expectation. With these in mind, we proceed to audit Google Search results to determine the extent to which government-maintained web domains, as authoritative sources on election information, are represented in the online environment of electoral information, during the 2022 US midterm elections. Specifically, we analyze where, how often, and how accurately state, county, and local government domains appear in Google’s organic results for election-related queries collected in 786 locations across the United States between October and November 2022, in a dataset of 3.45 million SERPs. Drawing parallels to studies of geotargeting errors in advertisements (Bandy and Hecht, 2021), we label the misalignment between the geographic area that a non-federal government web domain serves (such as a state or county) and the location for which it appears in search results as an instance of mistargeting, in which the search algorithm targets locality-specific information to a user in a different (incorrect) location. In this paper, we answer two related research questions:
@@
Findings and Contributions: We find that 40% of domains that appear among organic results are government websites, constituting nearly 40% of all organic results. However, in the 32.89% of organic results contributed by non-federal government domains, there is a 71.18% mistargeting rate. Our detailed analysis indicates a lack of consistency in the search algorithm’s treatment of locality-specific government sources, leaving most government domains, especially at the local level, appearing minimally, inconsistently, or missing entirely from election-related search results. We frame the omission and frequent mistargeting of non-federal government domains as a form of algorithmic misjudgement that obscures the important role that these domains play in providing official, locally-relevant information to voters. To the best of our knowledge, this is the largest locality-specific Google Search audit in the context of a political election.
@@ 3.1. Selection of Queries
A review of the query list after these steps revealed a lack of general queries about candidates and key voter issues, as well as a lack of biased and/or partisan queries. To fill these gaps, we augmented our query list by manually selecting queries formulated by eligible US voters included in the Voter Searches dataset in (Mustafaraj et al., 2020). Lastly, a few additional seed queries were added related to the topic of abortion, which was one of the top voting issues [^9] https://www.pewresearch.org/politics/2022/08/23/abortion-rises-in-importance-as-a-voting-issue-driven-by-democrats/ in the 2022 election. These processes resulted in a final list of 161 queries, of which 74 were from Google Trends. [^10] The query phrases and their provenance provenance information is available here: https://bit.ly/audit-queries.
Queries were reviewed qualitatively by one author to develop three major themes of a) Voting Logistics (in which queries are primarily about logistical information related to voting); b) Candidates and Issues (in which queries are primarily about candidates, key voting issues, and candidate’s opinions, stances, and/or leanings) and c) General Election Topics, (where queries are primarily general inquiries about the midterm election, such as what midterms are and when they occur). Three labelers then independently labeled all queries into these three categories. The majority label was used as the final label per query. Table 1 shows the query distribution by category and some examples for each category.
@@
Candidates and Issues who is on the ballot, candidate midterm info, candidates for my district 50
@@
Table 1. The audit used 161 queries that cover three major themes related to logistics of voting, candidates and key voting issues, and general inquiries about midterms. The query categories were assigned by three independent labelers.
@@ 3.2. Selection of Locations
We selected locations for data collection to correspond to U.S states and congressional districts whose House or Senate midterm races were classified as toss-ups, leaning Democrat, and leaning Republican, according to the Cook Political Report.111111 https://www.cookpolitical.com/ratings/senate-race-ratings and https://www.cookpolitical.com/ratings/house-race-ratings We identified 30 congressional districts with competitive House races across 23 states, as well as 10 states with competitive Senate races; 3 states had only competitive Senate races, making for a total of 26 unique states.
@@
To sample locations, we identified the median population of possible locations, and selected (up to) 15 locations below the median population and (up to) 15 locations above the median population, as well as the location with the highest population. Some states (for example, Nevada, which had only a competitive Senate race) had fewer than 31 cities, towns, and townships from the relevant counties. This process resulted in 872 locations across 26 states. Google Places API was used to get latitude and longitude coordinates for each of the locations. This information is needed to change the location of the Chrome browser to the desired location for the audit, [^15] Location information is available here: https://bit.ly/audit-locations as shown next.
@@ 3.3. Scraping Audit and SERP Validation
SERPs were collected for 30 days, starting on October 18, 2022 and ending November 18, 2022. Data were not collected on November 14 and 15. Data were collected using 12 desktop computers, with 2 to 3 states collecting per computer. Because of the large amount of locations, data collection started each day at 2am EST and continued until the evening. The order of locations was not changed daily, so SERPs per query were collected at approximately the same time within each location, and locations were completed sequentially. 3,838,128 SERPs were collected in total.
@@ 4. Methods
Since our research questions revolve around the presence of government-maintained websites in SERPs, we describe in this section how we extracted their occurrences; labeled them according to the government type; and defined when they are mistargeted with respect to the search location.
4.1. Extracting domains and election information boxes
SERPs are composed of multiple components that supplement the organic search results (the blue links) such as top stories, videos, images, the people also ask panel, tweets, and more. During election seasons, Google includes election-specific boxes on many SERPs with information about how to vote, notable dates, or interactive tools to find your voting location. Thus, we parsed SERPs for two types of information: a) the organic search results links; and b) the election information boxes.
Organic search links. 22,750 unique links were extracted from 31,310,263 organic results. Once all such links were extracted from all SERPs, they were cleaned to their root domain; however, given our focus on various government sources, in some cases we maintain the subdomain and avoid aggregating to the root, particularly if the subdomain is a county or town website. For example, while clintonco.illinois.gov is a subdomain of the state government website illinois.gov, we maintain the county government website. 4,556 unique domains were identified.
Election Information Boxes. We identified seven types of election boxes in the SERPs, which were displayed either at the top of the page (with a specific banner) or on the right side. The content of these boxes came from Democracy Works (for notable dates, where to vote, how to vote, etc.), Wikipedia (general information about the election), and Associated Press (election results).
@@ 4.2. Identifying Government Websites
Additionally, we identified some county and city websites using the .us domain that were not included in the GSA repository; because the .us domain is not strictly governmental, we manually reviewed all websites with the .us domain and included only those which corresponded to a county or city. 362 domains were added by manual review. In total, 1,848 websites were identified, which we call “government domains” or “government sources”. We are making the information about these domains in our dataset publicly available. [^18] https://bit.ly/FAccT24_gov_domains The distribution of domains across government types is shown in Table 2.
@@
Table 2. 1,848 government sources were identified from the data set of 4,556 domains (40.6%). Government sources were labeled for type and geoprovenance. We focus our analysis on state, county, and local government sources, as senator, representative, and Native sources show up minimally in organic results.
4.3. Quantifying Misjudgement
@@ 4.3.1. Domain level
Misjudgement can be studied to varying degrees of granularity. First, we examine misjudgement with respect to government domains and where they appear. To do so, we first identify how many government domains that appeared in organic results are from states in which we did not collect data; these are domains which are always mistargeted. Next, for domains originating in states for which we collected data, we calculate the proportion of times the domain appeared in the appropriate state and the proportion of times the domain was mistargeted (at the state-level). This gives insight into whether there are some locality-specific government domains that Google shows often across locations and are frequently mistargeted.
@@ 4.3.2. Aggregate Organic Results
We also study misjudgement by looking at all organic results in aggregate, and parsing differences in misjudgement by query categories, domain type, and state. Every organic result is labeled as to whether it is a government domain, and if so, we compare the domain’s geoprovenance (the geographic area it serves, such as a state, county, or city) to the search location. This gives us a more precise understanding of mistargeting than at the domain level, in which we only compare states. Organic results are labeled as either correct or mistargeted. Mistargeted results for county and city domains are also labeled as either in-state, if they serve a different location in the same state as the search location, or out-of-state, if they do not, which functions as an indicator for the severity of mistargeting. For each query category and state, as well as overall, we calculate the proportion of organic results composed of all non-federal government sources and the proportion of results composed of correctly targeted government results. Mistargeting rate is calculated by subtracting the proportion of correctly targeted from the total proportion, and dividing by the total proportion. We similarly calculate the proportion of results that are mistargeted but in-state, which gives insight into the nature of mistargeted results.
@@ 4.3.3. SERP Level
Lastly, we quantify misjudgement at the SERP level by calculating descriptive statistics about the number of government sources that appear per SERP, as well as descriptive statistics about the rank of correctly and incorrectly targeted government results. Since users heavily rely on ranking of search results (Pan et al., 2007), this informs the severity of misjudgement with respect to a user.
@@ 5. Results
Across all queries, local government sources are minimally represented among organic results. The prominence of state-level government sources compared to other government sources can be explained by the key roles that state-level actors play in elections: while the composition of election administrations vary by state, elections in each state are overseen by a chief election officer (such as a secretary of state, a governor, or a board), and their responsibilities include enforcing election laws, administering a statewide voter registration database, and other duties related to election logistics. [^19] https://www.ncsl.org/elections-and-campaigns/election-administration-at-state-and-local-levels However, their dominance is striking. While state government websites make up 3.5% of all websites (160 out of 4,556), they are shown in 30.06% of Voting Logistics results and in 19.67% of overall results.
On a SERP level, Voting Logistics queries had a median of 6 government results, with medians of 3 state-level and 2 county-level government sources, aligning with the aggregate proportions in Table3. In contrast, General Election Topics queries had a median of 2 government source per page and Candidates and Issues queries had a median of 1. The fact that the median of state-level websites per SERP is 3 for Voting Logistics queries naturally raises the question of whether all these three websites are targeted correctly to the search location.

### v2

@@ Abstract.
Google Search is an important way that people seek information about politics (Dutton et al., 2017), and Google states that it is “committed to providing timely and authoritative information on Google Search to help voters understand, navigate, and participate in democratic processes.” [^1] https://elections.google/civics-in-search/ This paper studies the extent to which government-maintained web domains are represented in the online electoral information environment, as captured through 3.45 Google Search result pages collected during the 2022 US midterm elections for 786 locations across the United States. Focusing on state, county, and local government domains that provide locality-specific information, we study not only the extent to which these sources appear in organic search results, but also the extent to which these sources are correctly targeted to their respective constituents. We label misalignment between the geographic area that non-federal domains serve and the locations for which they appear in search results as algorithmic mistargeting, a subtype of algorithmic misjudgement in which the search algorithm targets locality-specific information to users in different (incorrect) locations. In the context of the 2022 US midterm elections, we find that 71% of all occurrences of state, county, and local government sources were mistargeted, with some domains appearing disproportionately often among organic results despite providing locality-specific information that may not be relevant to all voters. However, we also find that mistargeting often occurs in low ranks. We conclude by considering the potential consequences of extensive mistargeting of non-federal government sources and argue that ensuring the correct targeting of these sources to their respective constituents is a critical part of Google’s role in facilitating access to authoritative and locally-relevant electoral information.
@@
journalyear: 2024copyright: rightsretainedconference: The 2024 ACM Conference on Fairness, Accountability, and Transparency; June 3–6, 2024; Rio de Janeiro, Brazilbooktitle: The 2024 ACM Conference on Fairness, Accountability, and Transparency (FAccT ’24), June 3–6, 2024, Rio de Janeiro, Brazildoi: 10.1145/3630106.3658916isbn: 979-8-4007-0450-5/24/06ccs: Information systems Web search enginesccs: Information systems Personalization
@@ 1. Introduction
Electoral denialism—the belief that elections are rigged or that electoral fraud is rampant—in the United States (and other countries, such as Brazil and Peru) is emerging as a significant threat to democracy (Cameron, 2023) and has roots in racial resentment and rising conspiracism (Stewart III, 2023). Since the early 2000s, internet was regarded as a medium “uniquely receptive to conspiracism” (James, 2000), and in the past two decades, there is little doubt that it has contributed significantly to conspiracism going mainstream (Rosenblum and Muirhead, 2019). Following the 2020 US Presidential Election, “election rigging” conspiracies that came to be labeled as “The Big Lie” led to the violent attack to the US Capitol on January 6, 2021 (Fahey, 2023). While beliefs about electoral fraud are somewhat persistent (Enders et al., 2021), the 2020 electoral denialism movement was different because it was embraced by partisan state legislatures, as reported by the Brennan Center for Justice. [^2] https://www.brennancenter.org/our-work/research-reports/voting-laws-roundup-june-2023 As a result, a large number of changes to electoral laws (some of them restricting and some of them expanding voting rights), became operational ahead of the 2022 US midterm election. [^3] https://www.reuters.com/legal/how-new-us-laws-could-trip-up-voters-midterm-elections-2022-11-01/
@@
Search engines also position themselves as trustworthy intermediaries for seeking election information. Google states that it is “committed to providing timely and authoritative information on Google Search to help voters understand, navigate, and participate in democratic processes.” [^6] https://elections.google/civics-in-search/ In an article about their work to “support” the 2022 US midterm elections, Google’s Vice President for Trust and Safety writes, “Our work is centered around connecting voters to the latest election information …” [^7] https://blog.google/outreach-initiatives/civics/our-ongoing-work-to-support-the-2022-us-midterm-elections/ and highlights “features that show data from nonpartisan organizations” that work to “connect voters with accurate information about voter registration and how to vote,” referring to Google Search election information panels that contain aggregated information from Democracy Works [^8] Democracy Works describes itself as “a nonpartisan, nonprofit organization that collaborates with election officials, leading tech platforms, and world-class partners to drive voter access and participation”. Source: https://www.democracy.works/about. about where and how to vote.
@@
Study Motivation: Given this background, three assumptions motivate the current study: (1) voting-eligible US citizens who search for electoral information using Google’s search engine have the right to receive accurate and relevant search results, regardless of where they live; (2) similar to Ballatore et al. (Andrea Ballatore and Sen, 2017), we acknowledge a spatial dimension to searching and search results, in which users searching for information related to elections and voting deserve results relevant to their geographic area, and 3) given the size of the United States and the varying voting laws per state (most of them updated ahead of the 2022 elections), receiving locality-specific information from authoritative sources is a reasonable expectation. Following these assumptions, we proceed to audit Google Search to determine the extent to which government-maintained web domains, as authoritative sources on election information, are represented in the online environment of electoral information, during the 2022 US midterm elections. Specifically, we analyze where, how often, and how accurately state, county, and local government domains appear in Google’s organic results for election-related queries collected in 786 locations across the United States between October and November 2022, in a dataset of 3.45 million search engine result pages (SERPs). Drawing parallels to studies of geotargeting errors in advertisements (Bandy and Hecht, 2021), we label the misalignment between the geographic area that a non-federal government web domain serves (such as a state or county) and the location for which it appears in search results as an instance of mistargeting, in which the search algorithm targets locality-specific information to a user in a different (incorrect) location. In this paper, we answer two related research questions:
@@
Findings and Contributions: We find that 40.6% (1,848/4,556) of domains that appear among organic results are government sources, and that these government sources constitute nearly 40% of all organic results occurrences. Despite this proportionality, we find that 71% of the occurrences of state, county, and local government sources were mistargeted, with a handful of locality-specific government domains appearing disproportionately often among organic results, despite providing locality-specific information that may not be relevant to all voters. Despite extensive mistargeting, we find that correctly targeted government sources usually appear in higher ranks, potentially mitigating possible harm inflicted by mistargeting. We conclude by considering how mistargeting factors into Google’s role in facilitating access to accurate, authoritative, and locally-relevant electoral information to voters. [^9] Audit and analysis materials, including query phrases, locations, and scripts, are available on GitHub: https://github.com/credlab/facct24
@@ 3.1. Selection of Queries
A review of the query list after these steps revealed a lack of general queries about candidates and key voter issues, as well as a lack of biased and/or partisan queries. To fill these gaps, we augmented our query list by manually selecting queries formulated by eligible US voters included in the Voter Searches dataset in (Mustafaraj et al., 2020). Lastly, a few additional seed queries were added related to the topic of abortion, which was one of the top voting issues in the 2022 election. [^10] https://www.pewresearch.org/politics/2022/08/23/abortion-rises-in-importance-as-a-voting-issue-driven-by-democrats/ These processes resulted in a final list of 161 queries, of which 74 were from Google Trends.
Queries were reviewed qualitatively by one author to develop three major themes of a) Voting Logistics (in which queries are primarily about logistical information related to voting); b) Candidates and Issues (in which queries are primarily about candidates and candidates’ opinions, stances, and/or leanings) [^11] The category Candidates and Issues includes 9 queries about abortion, a key voting issue in the 2022 US midterms. While these 9 queries could constitute their own category, we found that separating “Candidates” and “Issues” queries did not lead to meaningful changes in any results with respect to mistargeting rates, as government sources appear very minimally for these 9 queries. Thus, it was sufficient to lump these queries in with queries about candidates and candidate stances. Future work may study mistargeting for a more expansive set of queries representing different key voting issues. and c) General Election Topics, (where queries are primarily general inquiries about the midterm election, such as what midterms are and when they occur). Three labelers then independently labeled all queries into these three categories. The majority label was used as the final label per query. Table 1 shows the query distribution by category and some examples for each category.
@@
Candidates and Issues candidate midterm info, candidates for my district, candidate opinions 50
@@
Table 1. The audit used 161 queries that cover three major themes related to logistics of voting, candidates, and general inquiries about midterms. The query categories were assigned by three independent labelers.
@@ 3.2. Selection of Locations
We selected locations for data collection to correspond to U.S states and congressional districts whose House or Senate midterm races were classified as toss-ups, leaning Democrat, and leaning Republican, according to the Cook Political Report.121212 https://www.cookpolitical.com/ratings/senate-race-ratings and https://www.cookpolitical.com/ratings/house-race-ratings We identified 30 congressional districts with competitive House races across 23 states, as well as 10 states with competitive Senate races; 3 states had only competitive Senate races, making for a total of 26 unique states.
@@
To sample locations, we identified the median population of possible locations, and selected (up to) 15 locations below the median population and (up to) 15 locations above the median population, as well as the location with the highest population. Some states (for example, Nevada, which had only a competitive Senate race) had fewer than 31 cities, towns, and townships from the relevant counties. This process resulted in 872 locations across 26 states. Google Places API was used to get latitude and longitude coordinates for each of the locations. This information is needed to change the location of the Chrome browser to the desired location for the audit, as shown next.
@@ 3.3. Scraping Audit and SERP Validation
SERPs were collected for 30 days, starting on October 18, 2022 and ending November 18, 2022. Data were not collected on November 14 and 15. Data were collected using 12 desktop computers, with 2 to 3 states collecting per computer. Because of the large amount of locations, data collection started each day at 2am EST and continued until the evening. The order of locations was not changed daily, so SERPs per query were collected at approximately the same time within each location, and locations were completed sequentially. Due to random hardware or software failure, some SERPs failed to collect. 3,838,128 SERPs were collected in total (92.3% collection rate).
@@ 4. Methods
Since our research questions revolve around the presence of government-maintained websites in SERPs, we describe in this section how we extracted their occurrences, labeled them according to the government type and geographic area, and defined when they are mistargeted with respect to the search location.
4.1. Extracting domains
SERPs are composed of organic results (the blue links) and non-organic results that include top stories, videos, knowledge panels, and, during election season, election information panels. This study focuses on organic results only, and in particular the domains that occur in organic results. 22,750 unique links were parsed from 31,310,263 organic results. Once all such links were extracted from all SERPs, they were cleaned to their root domain; however, given our focus on various government sources, in some cases we maintain the subdomain and avoid aggregating to the root, particularly if the subdomain is a county or town website. For example, while clintonco.illinois.gov is a subdomain of the state government website illinois.gov, we maintain the county government website. 4,556 unique domains were identified. For each SERP, we additionally record the rank of the organic results. In determining rank, we do not account for the presence of non-organic panels that may be on the page; rank is based only on the organic results.
@@ 4.2. Identifying Government Websites
Additionally, we identified some county and city websites using the .us domain that were not included in the GSA repository; because the .us domain is not strictly governmental, we manually reviewed all websites with the .us domain and included only those which corresponded to a county or city. 362 domains were added by manual review. In total, 1,848 websites were identified, which we call “government domains” or “government sources’.’ The distribution of domains across government types is shown in Table 2.
@@
Table 2. 1,848 government sources were identified from the data set of 4,556 domains (40.6%). Government sources were labeled for type and geoprovenance. We focus our analysis on state, county, and local government sources, as senator, representative, and Native sources show up minimally, composing less than 0.1% of all organic results.
4.3. Measuring Mistargeting
We label misalignment between the geographic area that non-federal domains serve and the locations for which they appear in search results as mistargeting, a type of algorithmic misjudgement in which the search algorithm targets locality-specific information to users in different (incorrect) locations. Mistargeting can be studied at varying degrees of granularity.
@@ 4.3.1. Domain level
First, we examine mistargeting with respect to government domains and where they appear. To do so, we first identify how many government domains that appeared in organic results are from states in which we did not collect data; these are domains which are always mistargeted. Next, for domains originating in states for which we collected data, we calculate the proportion of times the domain appeared in the appropriate state and the proportion of times the domain was mistargeted (at the state-level). This gives insight into whether there are some locality-specific government domains that Google shows often across locations and are frequently mistargeted.
@@ 4.3.2. Aggregate Organic Results
We also study mistargeting by looking at all organic results in aggregate and and by parsing differences in mistargeting by query categories, domain type, and state. Every organic result is labeled as to whether it is a government domain, and if so, we compare the domain’s geoprovenance (the geographic area it serves, such as a state, county, or city) to the search location. This gives us a more precise understanding of mistargeting than at the domain level, in which we only compare states. Organic results are labeled as either correct or mistargeted. Mistargeted results for county and city domains are also labeled as either in-state, if they serve a different location in the same state as the search location, or out-of-state, if they do not, which functions as an indicator for the severity of mistargeting. For each query category, and across all queries, we calculate the proportion of organic results composed of all non-federal government sources and the proportion of results composed of correctly targeted government results. Mistargeting rate is calculated by subtracting the proportion of correctly targeted from the total proportion, and dividing by the total proportion. We similarly calculate the proportion of results that are mistargeted but in-state, which gives insight into the nature of mistargeted results.
@@ 4.3.3. SERP Level
Lastly, we quantify mistargeting at the SERP level by calculating descriptive statistics about the number of government sources that appear per SERP, as well as descriptive statistics about the rank of correctly and incorrectly targeted government results. Since users heavily rely on ranking of search results (Pan et al., 2007), this informs the severity of misjudgement with respect to a user.
@@ 5. Results
Across all queries, local government sources are minimally represented among organic results. The prominence of state-level government sources compared to other government sources may be explained by the key roles that state-level actors play in elections: while the composition of election administrations vary by state, elections in each state are overseen by a chief election officer (such as a secretary of state, a governor, or a board), and their responsibilities include enforcing election laws, administering a statewide voter registration database, and other duties related to election logistics. [^18] https://www.ncsl.org/elections-and-campaigns/election-administration-at-state-and-local-levels However, their dominance is striking. While state government websites make up 3.5% of all websites (160 out of 4,556), they compose 30.06% of Voting Logistics results and 19.67% of overall results.
On a SERP level, Voting Logistics queries had a median of 6 government results, with medians of 3 state-level and 2 county-level government sources, aligning with the aggregate proportions in Table3. In contrast, General Election Topics queries had a median of 2 government source per page and Candidates and Issues queries had a median of 1.